OpenStudy (anonymous):

simplify the expression: sqrt(27x^5) + x sqrt(3) + 9 sqrt(3x) If the simplified expression is written in standard form, what is the leading coefficient?

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

\[27x^5+x \sqrt{3}+9\sqrt{3x}\]is it yours?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no,\[\sqrt{27x^5}\] it is the first term, right?

OpenStudy (anonymous):

oh yea

OpenStudy (anonymous):

so, what do you want? simplify or find coefficient? haaa let electrokid helps you

OpenStudy (anonymous):

well it says to simplify but all i need is the leading coefficient

OpenStudy (anonymous):

\[\sqrt{27x^5}+x\sqrt{3}+9\sqrt{3x}\\ =3\sqrt{3}x^{5/3}+x\sqrt{3}+9\sqrt{3}\sqrt{x}\]

OpenStudy (anonymous):

to me, the first term = (27)^1/2 * x^5/2 the second one = sqr(3) *x yeap. agree

OpenStudy (anonymous):

okay, what's the leading coefficient?

OpenStudy (anonymous):

LOL do it. electrokid

OpenStudy (anonymous):

what is your "standard form"???

OpenStudy (anonymous):

the thing you told me, what's the coefficient of that?

OpenStudy (anonymous):

i believe, the leading term would be the term with the highest power of "x"

OpenStudy (anonymous):

which would be? I can't tell... :/

OpenStudy (anonymous):

what do you want? x is coefficient of number or vice versa?

OpenStudy (anonymous):

if that is what it is, then arrange the above expression in decresing order of powers of "x" from left to the right

OpenStudy (anonymous):

which is the leading coefficient of the simplified answer.

OpenStudy (anonymous):

umm ok

OpenStudy (anonymous):

the powers of x are: 5/3 1 1/2

OpenStudy (anonymous):

5/2

OpenStudy (anonymous):

@Hoa all yours

OpenStudy (anonymous):

5/3? isn't 1 bigger than 5/3? or would it be 3?

OpenStudy (anonymous):

@electrokid don't understand

OpenStudy (anonymous):

i think the leading coefficient of the standard form is 3.

OpenStudy (anonymous):

it is \(3\sqrt{3}\)

OpenStudy (anonymous):

my bad.. the powers of "x" are: 5/2 1 1/2

OpenStudy (anonymous):

okay thnx

OpenStudy (anonymous):

yer welcome.